# A modern perspective on rational homotopy theory

### v1

@@ 1. Introduction
Section 2 comprises background on $1$-reduced simplicial sets. In Section 3, we perform the left transfer to obtain Theorem 1.1. In Section 4, we prove some results about the fibrations in this left-induced model structure. In Section 5, we establish that the model category from Theorem 1.1 is simplicial. Section 6 consists of auxiliary characterization results from the rational homotopy theory of $1$-connected spaces. We use these results in Section 7 to perform the left Bousfield localization that yields Theorem 1.2. In Section 8, we give an alternative approach to Theorem 1.2 via Bousfield localization with respect to homology. In Section 9, we give an overview of rational homotopy theory beyond the setting of $1$-connectedness. In Section 10, we perform the left Bousfield localization that yields Theorem 1.3. In Section 11, we establish the analog of Theorem 1.3 for topological spaces. Lastly, in Section 12, we show that our localization of spaces with respect to rational homotopy is not equal to the Bousfield localization of spaces with respect to rational homology.
Acknowledgments
@@ 2. $1$-reduced simplicial sets
We begin with some background on $1$-reduced simplicial sets. We define the $1$-skeleton and the $1$-coskeleton of a simplicial set. Using these notions, we define $1$-reduced simplicial sets. Then, we define the $1$-reduction of a simplicial set, as well as the $1$st Eilenberg subcomplex of a pointed simplicial set. These constructions provide us with adjunctions between $1$-reduced simplicial sets and (pointed) simplicial sets that we rely on in this paper.
@@ 2.1. Simplicial sets
We collect background and notation on simplicial sets; a reference is [GJ99, Chapter I].
The combinatorics involved in the definition of a simplicial set can be packaged in the simplex category $\Delta$.
@@
The simplex category $\Delta$ is the category whose
  1. (1)
    objects are the sets $[n]=\{0,1,\dots,n\}$ for a non-negative integer $n$; and
  2. (2)
    morphisms are the order-preserving maps $\theta\colon[n]\rightarrow[m]$.
@@
For $0\leq i\leq n$, two special types of maps in $\Delta$ are
  1. (1)
    the $i$-th coface map $d^{i}\colon[n-1]\rightarrow[n]$; and
  2. (2)
    the $i$-th codegeneracy map $s^{i}\colon[n+1]\rightarrow[n]$.
These maps generate the simplex category $\Delta$ subject to a list of identities that determine their composites.
The simplex category $\Delta$ allows for a concise definition of a simplicial set. We write ${\mathcal{S}\mathrm{ets}}$ for the category of sets.
@@
A simplicial set is a functor $X\colon\Delta^{\operatorname{op}}\rightarrow{\mathcal{S}\mathrm{ets}}$. In detail, a simplicial set comprises
  1. (1)
    a set of $n$-simplices $X_{n}$ for every non-negative integer $n$;
  2. (2)
    face maps $d_{i}\colon X_{n}\rightarrow X_{n-1}$ for $0\leq i\leq n$; and
  3. (3)
    degeneracy maps $s_{i}\colon X_{n}\rightarrow X_{n+1}$ for $0\leq i\leq n$;
subject to a list of simplicial identities that govern how the faces and degeneracies are composed.
We write ${\mathcal{SS}\mathrm{ets}}$ for the category of simplicial sets and simplicial maps (natural transformations) between them.
@@
The standard $n$-simplex is the representable simplicial set
$\Delta[n]=\operatorname{Hom}_{\Delta}\left(-,[n]\right).$
The identity $\operatorname{id}_{[n]}$ gives an $n$-simplex $\sigma_{n}$ of $\Delta[n]$. For a simplicial set $X$, the Yoneda lemma yields the bijection
$X_{n}\cong\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}}\left(\Delta[n],X\right)$
that is natural in $[n]$ and in $X$. Therefore, we occasionally view an $n$-simplex of $X$ as a map $\Delta[n]\rightarrow X$.
Example 2.5.
The boundary $\partial\Delta[n]$ is the smallest subcomplex of $\Delta[n]$ containing the faces $d_{i}(\sigma_{n})$ for $0\leq i\leq n$. We write
$i_{n}\colon\partial\Delta[n]\rightarrow\Delta[n]$
for the boundary inclusion.
Example 2.6.
For $n\geq 1$ and $0\leq k\leq n$, the horn $\Lambda[n,k]$ is the smallest subcomplex of $\Delta[n]$ containing $d_{i}(\sigma_{n})$ for $i\neq k$. We write
$j_{n,k}\colon\Lambda[n,k]\rightarrow\Delta[n]$
for the horn inclusion. Observe that horns are undefined for $n=0$.
We can geometrically realize a simplicial set $X$ to obtain a CW complex $|X|$. Writing ${\mathcal{T}\mathrm{op}}$ for the category of topological spaces, we denote the geometric realization functor by
$|-|\colon{\mathcal{SS}\mathrm{ets}}\rightarrow{\mathcal{T}\mathrm{op}}.$
Conversely, for a topological space $Y$, its singular complex $\operatorname{Sing}(Y)$ is the simplicial set with $n$-simplices
$\operatorname{Sing}(Y)_{n}=\operatorname{Hom}_{\mathcal{T}\mathrm{op}}\left(|\Delta[n]|,Y\right).$
This pair of functors forms an adjunction that we use to oscillate between simplicial sets and topological spaces.
Proposition 2.7 ([GJ99, Chapter I, Proposition 2.2]).
The geometric realization functor $|-|\colon{\mathcal{SS}\mathrm{ets}}\rightarrow{\mathcal{T}\mathrm{op}}$ is left adjoint to the singular complex functor $\operatorname{Sing}\colon{\mathcal{T}\mathrm{op}}\rightarrow{\mathcal{SS}\mathrm{ets}}$.
Our work largely involves pointed simplicial sets, which are simplicial sets with a designated 00-simplex as their basepoint. We write ${\mathcal{SS}\mathrm{ets}}_{\ast}$ for the category of pointed simplicial sets and pointed simplicial maps between them. The following adjunction is verified just like its analog in topological spaces.
Lemma 2.8.
The forgetful functor $V\colon{\mathcal{SS}\mathrm{ets}}_{\ast}\rightarrow{\mathcal{SS}\mathrm{ets}}$ has a left adjoint
$(-)_{+}\colon{\mathcal{SS}\mathrm{ets}}\rightarrow{\mathcal{SS}\mathrm{ets}}_{\ast}$
defined by attaching a disjoint basepoint.
@@ 2.2. The $1$-skeleton and the $1$-coskeleton of a simplicial set
To define $1$-reduced simplicial sets, we recall the $1$-skeleton of a simplicial set. We also recall the $1$-coskeleton of a simplicial set, which we refer to in the definition of the $1$st Eilenberg subcomplex of a pointed simplicial set. Our reference is [Ber18, Chapter 2.6].
Let $\Delta_{1}$ denote the full subcategory of the simplex category $\Delta$ with objects $[0]$ and $[1]$. The inclusion
$\Delta_{1}\rightarrow\Delta$
induces a truncation functor
$\operatorname{tr}_{1}\colon{\mathcal{SS}\mathrm{ets}}\rightarrow{\mathcal{S}\mathrm{ets}}^{\Delta_{1}^{\operatorname{op}}}.$
This truncation functor admits both a left adjoint
$s_{1}\colon{\mathcal{S}\mathrm{ets}}^{\Delta_{1}^{\operatorname{op}}}\rightarrow{\mathcal{SS}\mathrm{ets}}$
and a right adjoint
$c_{1}\colon{\mathcal{S}\mathrm{ets}}^{\Delta_{1}^{\operatorname{op}}}\rightarrow{\mathcal{SS}\mathrm{ets}},$
in terms of which we can define the $1$-skeleton and the $1$-coskeleton of a simplicial set.
@@
Let $X$ be a simplicial set.
  1. (1)
    The $1$-skeleton of $X$ is the simplicial set
    $\operatorname{sk}_{1}(X)=s_{1}\operatorname{tr}_{1}(X).$
  2. (2)
    The $1$-coskeleton of $X$ is the simplicial set
    $\operatorname{cosk}_{1}(X)=c_{1}\operatorname{tr}_{1}(X).$
In detail, both $\operatorname{sk}_{1}(X)$ and $\operatorname{cosk}_{1}(X)$ have set of 00-simplices $X_{0}$ and set of $1$-simplices $X_{1}$, but they differ in their higher-dimensional simplices. The higher-dimensional simplices of $\operatorname{sk}_{1}(X)$ are the iterated degeneracies of its $1$-simplices. On the contrary, $\operatorname{cosk}_{1}(X)$ has an $n$-simplex for every compatible collection of $(n-1)$-simplices. Equivalently, $\operatorname{cosk}_{1}(X)$ has an $n$-simplex for every map $\operatorname{sk}_{1}(\Delta[n])\rightarrow X$.
@@
It follows from this definition that, for a simplicial set $X$, we have an inclusion map
$\operatorname{sk}_{1}(X)\rightarrow X$
and a quotient map
$X\rightarrow\operatorname{cosk}_{1}(X),$
where the latter map identifies the simplices of $X$ that share the same $1$-skeleton.
@@
2.3. $1$-reduced simplicial sets
We are now ready to define $1$-reduced simplicial sets.
@@ 2.3. $1$-reduced simplicial sets
A simplicial set $X$ is $1$-reduced if its $1$-skeleton $\operatorname{sk}_{1}(X)$ is a point. In detail, a simplicial set $X$ is $1$-reduced if it has a unique 00-simplex whose degeneracy is the unique $1$-simplex of $X$.
@@
Note that the geometric realization of a $1$-reduced simplicial set is a $1$-connected CW complex. It is also crucial to observe that a $1$-reduced simplicial set is unambiguously pointed, for it has a unique 00-simplex.
We write ${\mathcal{SS}\mathrm{ets}}_{1}$ for the category of $1$-reduced simplicial sets and simplicial maps between them.
@@
For $n\geq 2$, the simplicial $n$-sphere
$S_{\mathrm{simp}}^{n}=\Delta[n]/\partial\Delta[n]$
is a $1$-reduced simplicial set consisting of a unique 00-simplex and a unique non-degenerate $n$-simplex.
The proof of the following lemma is straightforward. We write $\wedge$ for the smash product.
@@ 2.4. The $1$-reduction of a simplicial set
We relate the category ${\mathcal{SS}\mathrm{ets}}_{1}$ of $1$-reduced simplicial sets to the category ${\mathcal{SS}\mathrm{ets}}$ of simplicial sets by means of a left adjoint to the forgetful functor $U\colon{\mathcal{SS}\mathrm{ets}}_{1}\rightarrow{\mathcal{SS}\mathrm{ets}}$. This left adjoint quotients out the $1$-skeleton of a simplicial set to yield a $1$-reduced simplicial set. Although quotients can be homotopically ill-behaved, this adjunction aids our model-categorical arguments later in this paper, for it repeatedly lets us translate lifting problems that we wish to solve into lifting problems that we can solve.
@@
The $1$-reduction of a simplicial set $X$ is the $1$-reduced simplicial set
$R_{1}(X)=X/\operatorname{sk}_{1}(X).$
@@
The $1$-reduction of a simplicial set is its largest $1$-reduced quotient in the sense of the following universal property, which is an instance of the universal property of quotients.
Lemma 2.18.
@@
Lemma 2.19.
The $1$-reduction ${R}_{1}\colon{\mathcal{SS}\mathrm{ets}}\rightarrow{\mathcal{SS}\mathrm{ets}}_{1}$ is left adjoint to the forgetful functor $U\colon{\mathcal{SS}\mathrm{ets}}_{1}\rightarrow{\mathcal{SS}\mathrm{ets}}$.
We note a consequence of this adjunction for later use.
@@
The claim is a consequence of Lemma 2.19 because $S_{\mathrm{simp}}^{2}=R_{1}(\Delta[2])$ by Example 2.17. ∎
2.5. The $1$st Eilenberg subcomplex of a pointed simplicial set
As we shall see, the approach of taking the $1$-reduction of a simplicial set does not play as nicely as we need it to for our homotopy-theoretic purposes, but we can do better. Because a $1$-reduced simplicial set has a unique 00-simplex, there is a forgetful functor
$F\colon{\mathcal{SS}\mathrm{ets}}_{1}\rightarrow{\mathcal{SS}\mathrm{ets}}_{\ast}$
from the category ${\mathcal{SS}\mathrm{ets}}_{1}$ of $1$-reduced simplicial sets to the category ${\mathcal{SS}\mathrm{ets}}_{\ast}$ of pointed simplicial sets. This functor has not only a left adjoint given by $1$-reduction, but also a right adjoint given by the $1$st Eilenberg subcomplex of a pointed simplicial set. We present its definition, though we do not use it beyond this section, as its only feature that we need in this paper is the aforementioned adjunction, which we establish.
Recall from Remark 2.10 that, for a simplicial set $X$, we have a quotient map $X\rightarrow\operatorname{cosk}_{1}(X)$.
Definition 2.21 ([May67, Definition 8.3]).
The $1$st Eilenberg subcomplex $E_{1}(X)$ of a pointed simplicial set $X$ with basepoint $x_{0}$ is defined by the pullback square
${{E_{1}(X)}}$${X}$${{\Delta[0]}}$${{\operatorname{cosk}_{1}(X).}}$$\scriptstyle{x_{0}}$
In detail, $E_{1}(X)$ consists of all simplices $\chi$ of $X$ such that every 00-dimensional face of $\chi$ is $x_{0}$ and every $1$-dimensional face of $\chi$ is the degeneracy of $x_{0}$.
It follows from this definition that the $1$st Eilenberg subcomplex $E_{1}(X)$ of a pointed simplicial set $X$ is $1$-reduced.
The $1$st Eilenberg subcomplex of a pointed simplicial set is its largest $1$-reduced subcomplex in the sense of the following universal property that is straightforward to verify.
Lemma 2.22.
Every pointed map $f\colon X\rightarrow Y$ from a $1$-reduced simplicial set $X$ to a pointed simplicial set $Y$ maps into the $1$st Eilenberg subcomplex $E_{1}(Y)$ of $Y$. Thus, there is a unique map $\hat{f}\colon X\rightarrow E_{1}(Y)$ such that the diagram
${X}$${{E_{1}(Y)}}$${{Y}}$$\scriptstyle{\hat{f}}$$\scriptstyle{f}$

### v2

@@ 1.4. Organization of the paper
Section 2 comprises background on $1$-reduced simplicial sets. In Section 3, we perform the left transfer. In Section 4, we investigate the fibrations in our left-induced model category. In Section 5, we establish that our left-induced model category is simplicial. Section 6 consists of auxiliary characterization results from the rational homotopy theory of $1$-connected spaces. We use these results in Section 7 to perform the left Bousfield localization that recovers Quillen’s family of model structures on $1$-reduced simplicial sets. In Section 8, we give an alternative approach via Bousfield localization with respect to homology. In Section 9, we give an overview of the rational homotopy theory of non-simply connected spaces. Lastly, in Section 10, we perform our left Bousfield localization to all spaces to generalize Quillen’s family of model structures.
1.5. Acknowledgments
@@ 2. $1$-reduced simplicial sets
We begin with some background on $1$-reduced simplicial sets. We first give background on simplicial sets; in particular, we define the $1$-skeleton and the $1$-coskeleton of a simplicial set. Using these notions, we define $1$-reduced simplicial sets. Then, we define the $1$-reduction of a simplicial set, as well as the 11st Eilenberg subcomplex of a pointed simplicial set. These constructions provide us with adjunctions between $1$-reduced simplicial sets and (pointed) simplicial sets that we rely on to bridge our investigation of the homotopy theory of $1$-reduced simplicial sets to known results from the homotopy theory of (pointed) simplicial sets.
@@ 2.1. Simplicial sets
We collect some background and notation on simplicial sets; a reference is [23, Chapter I]. The combinatorics involved in the definition of a simplicial set can be packaged in the simplex category $\Delta$.
@@
The simplex category $\Delta$ is the category whose objects are the sets $n]=\{0,1,\dots,n\}$ for a non-negative integer $n$ and whose morphisms are the order-preserving maps of these sets.
The order-preserving maps of finite total orders can be built out of special injections and surjections.
@@
For $0\leq i\leq n$, two special maps in $\Delta$ are the $i$-th coface map $d^{i}\colon[n-1]\rightarrow[n]$, given by
$d^{i}(x)=\begin{cases}x,&x<i;\\ x+1,&x\geq i,\end{cases}$
and the $i$-th codegeneracy map $s^{i}\colon[n+1]\rightarrow[n]$, given by
$s^{i}(x)=\begin{cases}x,&x\leq i;\\ x-1,&x>i.\end{cases}$
These maps generate $\Delta$ subject to a list of identities on their composites.
The simplex category $\Delta$ yields a concise definition of a simplicial set; we write $\mathcal{S}\mathrm{ets}$} for the category of sets.
@@
A simplicial set is a functor $X\colon\Delta^{\operatorname{op}}\rightarrow{\mathcal{S}\mathrm{ets}}$. In detail, a simplicial set comprises a set of $n$-simplices $X_{n}$ for every non-negative integer $n$, face maps $d_{i}\colon X_{n}\rightarrow X_{n-1}$ for $0\leq i\leq n$, and degeneracy maps $s_{i}\colon X_{n}\rightarrow X_{n+1}$ for $0\leq i\leq n$ subject to a list of identities on how the faces and degeneracies are composed.
We write $\mathcal{SS}\mathrm{ets}$} for the category of simplicial sets and their simplicial maps (natural transformations). We give the fundamental building block examples of simplicial sets, which are reminiscent of simplicial complexes.
@@
The standard $n$-simplex is $\Delta[n]=\operatorname{Hom}_{\Delta}(-,[n]);$ the identity $\operatorname{id}_{[n]}$ gives an $n$-simplex $\sigma_{n}$. The boundary $\partial\Delta[n]$ is the smallest subcomplex of $\Delta[n]$ containing $d_{i}(\sigma_{n})$ for $0\leq i\leq n$. For $n\geq 1$ and $0\leq k\leq n$, the horn $\Lambda[n,k]$ is the smallest subcomplex of $\Delta[n]$ containing $d_{i}(\sigma_{n})$ for $i\neq k$.
As is the case for simplicial complexes, we can geometrically realize a simplicial set $X$ to obtain a CW complex $|X|$. Writing $\mathcal{T}\mathrm{op}$} for the category of topological spaces, we denote the geometric realization functor by $|-|\colon{\mathcal{SS}\mathrm{ets}}\rightarrow{\mathcal{T}\mathrm{op}}$. Its right adjoint sends a topological space $Y$ to the singular complex $\operatorname{Sing}(Y)$ with $n$-simplices $\operatorname{Sing}(Y)_{n}=\operatorname{Hom}_{\mathcal{T}\mathrm{op}}(|\Delta[n]|,Y)$. As the name suggests, there is a connection between the singular complex of a space and its singular homology; the $n$-th abelian group in the singular chain complex of a space is defined to be the free abelian group on the set of $n$-simplices of the singular complex of that space.
@@ 2.2. The $1$-skeleton and the $1$-coskeleton of a simplicial set
To define $1$-reduced simplicial sets, we recall the $1$-skeleton of a simplicial set. We also recall the $1$-coskeleton of a simplicial set, which we refer to in the definition of the 11st Eilenberg subcomplex of a pointed simplicial set. Our reference is [5, Chapter 2.6].
Let $\Delta_{1}$ denote the full subcategory of $\Delta$ with objects $[0]$ and $[1]$. The inclusion $\Delta_{1}\rightarrow\Delta$ induces a truncation functor $\operatorname{tr}_{1}\colon{\mathcal{SS}\mathrm{ets}}\rightarrow{\mathcal{S}\mathrm{ets}}^{\Delta_{1}^{\operatorname{op}}}$ that admits both a left adjoint $s_{1}\colon{\mathcal{S}\mathrm{ets}}^{\Delta_{1}^{\operatorname{op}}}\rightarrow{\mathcal{SS}\mathrm{ets}}$ and a right adjoint $c_{1}\colon{\mathcal{S}\mathrm{ets}}^{\Delta_{1}^{\operatorname{op}}}\rightarrow{\mathcal{SS}\mathrm{ets}}$, in terms of which we can define the $1$-skeleton and the $1$-coskeleton of a simplicial set.
@@
Let $X$ be a simplicial set. The $1$-skeleton of $X$ is $\operatorname{sk}_{1}(X)=s_{1}\operatorname{tr}_{1}(X)$. The $1$-coskeleton of $X$ is $\operatorname{cosk}_{1}(X)=c_{1}\operatorname{tr}_{1}(X)$.
In detail, both $\operatorname{sk}_{1}(X)$ and $\operatorname{cosk}_{1}(X)$ have set of 0-simplices $X_{0}$ and set of $1$-simplices $X_{1}$, but they differ in their higher-dimensional simplices. The higher-dimensional simplices of $\operatorname{sk}_{1}(X)$ are the iterated degeneracies of its $1$-simplices. On the contrary, $\operatorname{cosk}_{1}(X)$ has an $n$-simplex for every compatible collection of (n1)(n-1)-simplices. More precisely, $\operatorname{cosk}_{1}(X)$ has an $n$-simplex for every map $\operatorname{sk}_{1}(\Delta[n])\rightarrow X$.
@@
It follows that, for a simplicial set $X$, we have an inclusion map $\operatorname{sk}_{1}(X)\rightarrow X$ and a canonical map $X\rightarrow\operatorname{cosk}_{1}(X)$, where the latter map identifies the simplices of $X$ that share the same $1$-skeleton.
Besides the above remark, we also need the following example for later use in this section.
@@
2.3. Definition of $1$-reduced simplicial sets
We are now ready to introduce $1$-reduced simplicial sets.
@@ 2.3. Definition of $1$-reduced simplicial sets
A simplicial set $X$ is $1$-reduced if its $1$-skeleton $\operatorname{sk}_{1}(X)$ is a point. In detail, a simplicial set $X$ is $1$-reduced if it has a unique 0-simplex whose degeneracy is the unique $1$-simplex of $X$.
@@
Note that the geometric realization of a $1$-reduced simplicial set is a $1$-connected CW complex. It is also crucial to observe that a $1$-reduced simplicial set is unambiguously pointed, for it has a unique 0-simplex.
We write ${\mathcal{SS}\mathrm{ets}}_{1}$ for the category of $1$-reduced simplicial sets and simplicial maps between them. The most important example of a $1$-reduced simplicial set for our work is the following.
@@
For $n\geq 2$, the simplicial $n$-sphere $S_{\mathrm{simp}}^{n}=\Delta[n]/\partial\Delta[n]$ is a $1$-reduced simplicial set consisting of a unique 0-simplex and a unique non-degenerate $n$-simplex.
We also need the following lemma, whose proof is straightforward; we write $\wedge$ to denote the smash product.
@@ 2.4. The $1$-reduction of a simplicial set
We relate the category ${\mathcal{SS}\mathrm{ets}}_{1}$ of $1$-reduced simplicial sets to the category $\mathcal{SS}\mathrm{ets}$} of simplicial sets by means of a left adjoint to the forgetful functor $U\colon{\mathcal{SS}\mathrm{ets}}_{1}\rightarrow{\mathcal{SS}\mathrm{ets}}$. This left adjoint quotients out the $1$-skeleton of a simplicial set to yield a $1$-reduced simplicial set. Although quotients can be homotopically ill-behaved, this adjunction aids our model-categorical arguments in this paper, for it lets us translate lifting problems that we wish to solve in ${\mathcal{SS}\mathrm{ets}}_{1}$ into lifting problems that we can solve in $\mathcal{SS}\mathrm{ets}$}.
@@
The $1$-reduction of a simplicial set $X$ is the $1$-reduced simplicial set $R_{1}(X)=X/\operatorname{sk}_{1}(X)$.
The following observation carries useful consequences for later, so we formulate it explicitly.
@@
The $1$-reduction of a simplicial set is its largest $1$-reduced quotient in the sense of the following universal property, which is a consequence of the universal property of quotients.
Proposition 2.14.
@@
Proposition 2.15.
The $1$-reduction ${R}_{1}\colon{\mathcal{SS}\mathrm{ets}}\rightarrow{\mathcal{SS}\mathrm{ets}}_{1}$ is left adjoint to the forgetful functor.
We invoke Example 2.13 to note a consequence of this adjunction for later use.
@@
The claim is a consequence of Proposition 2.15 because $S_{\mathrm{simp}}^{2}=R_{1}(\Delta[2])$ by Example 2.13. ∎
2.5. The 11st Eilenberg subcomplex of a pointed simplicial set
As we shall see, the approach of taking the $1$-reduction of a simplicial set does not play as nicely as we need it to for our homotopy-theoretic purposes, but we can do better. Because a $1$-reduced simplicial set has a unique 0-simplex, there is a forgetful functor $F\colon{\mathcal{SS}\mathrm{ets}}_{1}\rightarrow{\mathcal{SS}\mathrm{ets}}_{\ast}$ from the category ${\mathcal{SS}\mathrm{ets}}_{1}$ of $1$-reduced simplicial sets to the category $\mathcal{SS}\mathrm{ets}$$_{\ast}$ of pointed simplicial sets. This functor also has a right adjoint given by the 11st Eilenberg subcomplex of a pointed simplicial set, which is its largest $1$-reduced subcomplex containing the basepoint.
Note that, in the absence of a basepoint, there is generally no canonical choice of such a subcomplex, so the use of pointed simplicial sets is essential for this construction. In fact, we shall see that the forgetful functor from $1$-reduced simplicial sets to simplicial sets has no right adjoint.
We present the definition of the 11st Eilenberg subcomplex, though the only feature that we need in this paper is the adjunction that it features in. Recall from Remark 2.6 that, for a simplicial set $X$, we have a canonical map $X\rightarrow\operatorname{cosk}_{1}(X)$ to its $1$-coskeleton.
Definition 2.17 ([30, Definition 8.3]).
The 11st Eilenberg subcomplex $E_{1}(X)$ of a pointed simplicial set $X$ with basepoint $x_{0}$ is defined by the pullback
[…]
In detail, $E_{1}(X)$ consists of all simplices $\chi$ of $X$ such that every 0-dimensional face of $\chi$ is $x_{0}$ and every $1$-dimensional face of $\chi$ is the degeneracy of $x_{0}$.
It follows that the 11st Eilenberg subcomplex $E_{1}(X)$ of a pointed simplicial set $X$ is $1$-reduced. The 11st Eilenberg subcomplex of a pointed simplicial set is its largest $1$-reduced subcomplex in the sense of the following universal property that is straightforward to verify.
Proposition 2.18.
Every pointed map $f\colon X\rightarrow Y$ from a $1$-reduced simplicial set $X$ to a pointed simplicial set $Y$ maps into the 11st Eilenberg subcomplex $E_{1}(Y)$ of $Y$. Thus, there is a unique map $\hat{f}\colon X\rightarrow E_{1}(Y)$ such that the diagram
[…]